# Adaptive Regularization for Large-Scale Sparse Feature Embedding Models

### v1

@@ Abstract
The one-epoch overfitting problem has drawn widespread attention, especially in CTR and CVR estimation models in search, advertising, and recommendation domains. These models which rely heavily on large-scale sparse categorical features, often suffer a significant decline in performance when trained for multiple epochs. Although recent studies have proposed heuristic solutions, they have not clearly identified the fundamental cause of this phenomenon. In this work, we provide a theoretical analysis that explains why overfitting occurs in models that use large-scale sparse categorical features. Based on this analysis, we propose an adaptive regularization method to address it. Our approach not only prevents the severe performance degradation observed during multi-epoch training, but also improves model performance within a single epoch. This method has already been deployed in online production systems.
@@ 1 Introduction
Click-through rate (CTR) and conversion rate (CVR) estimation are critical for advertising, search and recommendation (ASR) applications. E-commerce platforms like Amazon and Taobao rely on optimizing CTR and CVR estimation to boost gross merchandise volume (GMV), while advertising platforms at Google and Meta depend on it to drive revenue growth. In the past decade, as deep learning has been widely adopted in ASR applications, most estimation models have been built on deep learning frameworks and rely on large-scale, sparse categorical features (cheng2016wide; guo2017deepfm; lian2018xdeepfm). Recent work (zhang2022towards) demonstrates, through extensive experiments, that such models commonly suffer from the one-epoch overfitting phenomenon, where model performance drops sharply after the first epoch of training. It empirically suggests that the optimal practice is to train the models for only one epoch. At the same time, their empirical analysis also reveals a strong connection between feature sparsity and the one-epoch phenomenon. In ASR scenarios, categorical features can easily reach the scale of billions, and most feature values occur extremely infrequently (xie2020kraken; zhang2022picasso), which makes models especially susceptible to overfitting after the first epoch. Additionally, ouyang2022training reported a similar one-epoch phenomenon during supervised fine-tuning (SFT) of large language model (LLM), though they argue that moderate overfitting can actually be beneficial. We leave it in future work. In this paper, we focus on addressing the multi-epoch overfitting problem for models with large-scale sparse features in ASR applications.
@@
liu2023multi was the first to address the one-epoch problem using a heuristic approach. They introduced a multi-epoch data augmentation method (MEDA) that can be easily applied to estimation models with large-scale sparse categorical features. Later, they extended it to a continual learning paradigm (fan2024multi). To mitigate overfitting during multi-epoch training, MEDA reinitializes the embeddings of categorical features and their corresponding optimizer states at the beginning of each epoch. Although this approach effectively alleviates multi-epoch overfitting, it is fundamentally heuristic because the embedding parameters are only reinitialized at epoch boundaries. It neither ensures optimal convergence nor explains the root cause of overfitting. Moreover, this embedding reinitialization approach may result in the loss of substantial information, potentially resulting in suboptimal model performance. Nevertheless, it provides important inspiration for subsequent research. Another method, proposed by wang2025scaling, adopts an LLM-style paradigm. In this approach, generative pretraining is used to obtain embeddings for categorical features, which remain frozen during subsequent training. This strategy also helps to mitigate overfitting. However, it requires significant training resources to obtain the pretrained embeddings, and the performance cannot be fairly compared to single-epoch training because the parameter budget used during the pretraining phase is not included. Therefore, it cannot be regarded as a solution to the one-epoch problem, although it does highlight the pivotal role of the embedding layer in this phenomenon.
@@
There are several well-established and general-purpose methods to mitigate overfitting, such as dropout (srivastava2014dropout), $L_{1}$ and $L_{2}$ regularization (schmidt2009optimization; moore2011l1), and weight decay (hanson1988comparing). However, in real-world industrial applications, data is typically high-dimensional and sparse, with different features exhibiting varying degrees of sparsity. Consequently, methods like AdamW (loshchilov2017decoupled), which apply the same weight decay across all parameters, tend to be suboptimal for addressing overfitting. Specifically, such approaches may reduce the fitting accuracy for relatively dense features while failing to effectively control overfitting in sparse features.
@@ 2 Preliminary
In the industrial ASR domain, mainstream estimation models typically combine embedding layers with a variety of dense multi-layer perceptron (MLP) backbones, such as gu2022wukong. Taking an e-commerce platform estimation model as an example, these models process large-scale sparse inputs, including item IDs, brand IDs, and seller IDs, where the scale of categorical features can range from millions to billions. Before being processed by the estimation model, these sparse features are first mapped to low-dimensional dense representations via embedding layers. The one-hot categorical feature ${\bm{x}}_{i}(t)\in\mathbb{R}^{N_{i}}$ is defined as ${\bm{x}}_{i}(t)=\begin{bmatrix}0,0,\ldots,1,\ldots,0\end{bmatrix}^{\top},i\in[S],t\in[T]$, which can be represented by an embedding vector ${\bm{e}}_{i}\left(t\right)\in\mathbb{R}^{d_{i}}$ shown below
@@ 2.1 Neural Network Definition
where $L$ denotes the number of MLP layers, $\sigma(\cdot)$ is the ReLU function (agarap2018deep), ${\bm{W}}_{l}$ is the linear projection matrix of layer $l,l\in[L]$. In this paper, we only discuss the impact of embedding layer so we can define the $t$-th sample output of DNN as
@@ 2.2 Rademacher Complexity Bound
Rademacher complexity is a standard tool to control the uniform convergence of given classes of predictors (koltchinskii2002empirical,zhang2016understanding). Formally, given a real-valued function class $\mathcal{H}$ with the input $\bm{z}$}, we define the empirical Rademacher complexity $\widehat{\mathcal{R}}_{T}\left(\mathcal{H}\right)$ as
@@
Based on Theorem 1 of golowich2018size, and noting that the embedding layer can be regarded as a linear projection, we can obtain the following upper bound on the Rademacher complexity of equation 3 as
@@
where $\|\cdot\|$ denotes $\ell_{2}$ norm, and each matrix ${\bm{W}}_{l}$ in function $f$ has Frobenius norm at most $M_{F}(l),l\in[L]$. $M_{E_{i}}$ is the Frobenius norm of embedding matrix ${\bm{E}}_{i}$ and the activation functions are assumed to be 1-Lipschitz and positive-homogeneous. Furthermore, this upper bound can be rewritten as
$\widehat{\mathcal{R}}_{T}(\mathcal{H}_{L})\leq\frac{\left(\prod_{l=1}^{L}M_{F}(l)\right)\sqrt{S}}{\sqrt{T}}\left(\sqrt{\sum_{i=1}^{S}\sum_{j=1}^{N_{i}}\tau_{ij}}\right)\left(\sqrt{2\log(2)L}+1\right)$ (6)
where $\tau_{ij}$ is the squared $\ell_{2}$ norm of $j$-th row of embedding matrix ${\bm{E}}_{i}$. In ASR applications, the majority of parameters are concentrated in the embedding layers. Thus, we can see that $\sum_{i=1}^{S}\sum_{j=1}^{N_{i}}\tau_{ij}$ has a significant impact on the upper bound of Rademacher complexity, and consequently, on the generalization error bound.
@@ 3 The Proposed Approach
where $\varphi(\tau_{ij})=\min_{\|{\bm{e}}_{ij}\|^{2}\leq\tau_{ij}}\mathcal{L}({\bm{e}}_{ij})$ with $i\in[S]$ and $j\in[N_{i}]$. Here, ${\bm{e}}_{ij}$ denotes the $j$-th row of embedding matrix ${\bm{E}}_{i}$, and $\mathcal{L}({\bm{e}}_{ij})$ represents the average cross-entropy (CE) loss evaluated on the DNN output when the training sample activates ${\bm{e}}_{ij}$. For the CE loss function, $\mathcal{L}({\bm{e}}_{ij})$ is lower semicontinuous and bounded below on bounded sets (goodfellow2016deep). Consequently, $\varphi(\tau_{ij})$ is well-defined (ok2011real) and monotonically non-increasing in $\tau_{ij}$, since the feasible set expands with $\tau_{ij}$. Here, $\tau_{ij}$ is the squared norm budget of embedding vector ${\bm{e}}_{ij}$. According to dibenedetto2016real, $\varphi(\tau_{ij})$ is differentiable almost everywhere, and we assume that $\varphi(\tau_{ij})$ is differentiable at the optimal points $\tau^{*}_{ij}$ for analysis. The coefficient $m_{ij}$ denotes the sample frequency with which embedding vector ${\bm{e}}_{ij}$ appears in the training dataset and $C$ is a constant that defines the global norm upper bound for embedding layers. As shown in equation 6, the value of $C$ directly determines the Rademacher complexity upper bound.
@@
Proof. Based on the above assumption, at points of differentiability where a standard constraint qualification holds for the inner optimal solution, the envelope theorem and Lagrangian decomposition (shapiro1979survey) yield
@@
Based on the KKT condition (boyd2004convex), we have $m_{ij}\varphi^{\prime}(\tau^{*}_{ij})+\mu_{0}-\mu_{ij}=0$. Applying the complementary slackness condition, we obtain $\mu_{ij}\tau^{*}_{ij}=0$. This implies the optimality condition for $\tau_{ij}>0$ can be simplified as
@@ 3.1 Adaptive Regularization Method
As shown in equation 6, embedding layers in ASR applications typically take the majority of the model parameters and have a substantial impact on the generalization error bound. Proposition 1 suggests that the norm budget for each embedding vector should be allocated according to its sample frequency. However, it is not easy to use the frequency directly during the training process. We can estimate $m_{ij}$ via the stochastic occurrence interval $I_{ij}$ of ${\bm{e}}_{ij}$ for $i\in[S]$ and $j\in[N_{i}]$. Specifically, given that ${\bm{x}}_{i}(t)$ is sampled from an i.i.d. distribution (a common assumption in deep learning), we have $\mathbb{E}[m_{ij}]=T/\mathbb{E}[I_{ij}]$ (papoulis1965random). It enables us to incorporate frequency estimation into the norm budget allocation strategy during training.
Based on the analysis above, we propose an adaptive method that assigns the regularization strength based on the occurrence interval of each embedding vector. Specifically, at each training step $k$, we define last valid update step (LVS) for the embedding vector ${\bm{e}}_{ij}$ as $s_{ij}$. If the gradient norm of ${\bm{e}}_{ij}$ satisfies $\|{\bm{g}}_{ij}\|>0$, we update the LVS by setting $s_{ij}=k$. Otherwise, $s_{ij}$ retains its previous value. Therefore, $s_{ij}$ serves as a lazy-update variable. We define the stochastic update interval as $I_{ij}=k-s_{ij}-1$. The feature-wise regularization coefficient $\lambda_{ij}$ is then dynamically computed as below
$\lambda_{ij}=\min\left(1,\alpha I_{ij}\right),i\in[S],j\in[N_{i}]$ (11)
Here, $\alpha\in[0,1)$ denotes the base regularization coefficient. Following the decoupled weight decay approach in AdamW (loshchilov2017decoupled), we incorporate the dynamically computed regularization into each optimizer update step. Algorithm 1 outlines the procedure of Adam with adaptive regularization (AdamAR).
@@
In fact, our method is not only suitable for Adam, but also compatible with various gradient-based optimizers which have weight decay factor, such as Adagrad (duchi2011adaptive). We have implemented an adaptive version for Adagrad in algorithm 2 of the appendix B. In section 4, we conducted comparative experiments evaluating performance across both Adam and Adagrad.
@@ 3.2 Discussion on the Mechanism of Regularization
When adaptive regularization is applied according to equation 11, the update rule for parameters satisfy $\|\theta^{k-1}_{p}-\lambda^{t}_{p}\theta^{k-1}_{p}-\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}}+\varepsilon)\|\leq\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|+\|\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}}+\varepsilon)\|$
@@
$\|\theta^{k-1}_{p}-\lambda^{k}_{p}\theta^{k-1}_{p}-\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}}+\varepsilon)\|\leq\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|+\|\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}}+\varepsilon)\|$ (13)
From proposition 2, the update rule for $\theta^{k}_{p}$ can be interpreted intuitively. If the interval $I^{k}_{p}$ is large for the corresponding sparse categorical features, the term $\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|$ becomes negligible, and $\theta^{k}_{p}$ is effectively determined by the latest gradient. In the ASR domain, certain categorical feature values may not appear in every mini-batch, resulting in their embedding parameters being updated infrequently, while the MLP parameters are updated in every batch. After the MLP component has nearly converged, the embedding parameters associated with sparse features will have received far fewer updates and may become misaligned with the current state of the MLP parameters, making the previous value $\theta^{k-1}_{p}$ less informative. Proposition 2 demonstrates that the factor $\left(1-\alpha\right)^{I^{k}_{p}}$ exponentially attenuates the previous value $\theta^{k-1}_{p}$, making the embedding parameters of sparse features depend more heavily on the current gradient. Meanwhile, since the update interval $I^{k}_{p}$ for MLP parameters is identically zero (as they are updated in every batch), the regularization primarily affects the embedding parameters corresponding to very low-frequency feature values.
Moreover, we observe that the method proposed in liu2023multi is a special case of our approach when zero reinitialization is applied to the embedding and $I^{k}_{p}$ is specified as

### v2

@@ Abstract
The one-epoch overfitting problem has drawn widespread attention, especially in CTR and CVR estimation models in search, advertising, and recommendation domains. These models which rely heavily on large-scale sparse categorical features, often suffer a significant decline in performance when trained for multiple epochs. Although recent studies have proposed heuristic solutions, the fundamental cause of this phenomenon remains unclear. In this work, we present a theoretical explanation grounded in Rademacher complexity, supported by empirical experiments, to explain why overfitting occurs in models with large-scale sparse categorical features. Based on this analysis, we propose a regularization method that constrains the norm budget of embedding layers adaptively. Our approach not only prevents the severe performance degradation observed during multi-epoch training, but also improves model performance within a single epoch. This method has already been deployed in online production systems.
@@ 1 Introduction
Click-through rate (CTR) and conversion rate (CVR) estimation are critical for advertising, search and recommendation (ASR) applications. E-commerce platforms like Amazon and Taobao rely on optimizing CTR and CVR estimation to boost gross merchandise volume (GMV), while advertising platforms at Google and Meta depend on it to drive revenue growth. In the past decade, as deep learning has been widely adopted in ASR applications, most estimation models have been built on deep learning frameworks and rely on large-scale, sparse categorical features (Cheng et al. (2016); Guo et al. (2017); Lian et al. (2018)). Recent work (Zhang et al. (2022b)) demonstrates, through extensive experiments, that such models commonly suffer from the one-epoch overfitting phenomenon, where model performance drops sharply after the first epoch of training. It empirically suggests that the optimal practice is to train the models for only one epoch. At the same time, their empirical analysis also reveals a strong connection between feature sparsity and the one-epoch phenomenon. In ASR scenarios, categorical features can easily reach the scale of billions, and most feature values occur extremely infrequently (Xie et al. (2020); Zhang et al. (2022a)), which makes models especially susceptible to overfitting after the first epoch. Additionally, Ouyang et al. (2022) reported a similar one-epoch phenomenon during supervised fine-tuning (SFT) of large language model (LLM), though they argue that moderate overfitting can actually be beneficial. We leave it in future work. In this paper, we focus on addressing the multi-epoch overfitting problem for models with large-scale sparse features in ASR applications.
@@
Liu et al. (2023) was the first to address the one-epoch problem using a heuristic approach. They introduced a multi-epoch data augmentation method (MEDA) that can be easily applied to estimation models with large-scale sparse categorical features. Later, they extended it to a continual learning paradigm (Fan et al. (2024)). To mitigate overfitting during multi-epoch training, MEDA reinitializes the embeddings of categorical features and their corresponding optimizer states at the beginning of each epoch. Although this approach effectively alleviates multi-epoch overfitting, it is fundamentally heuristic because the embedding parameters are only reinitialized at epoch boundaries. It neither ensures optimal convergence nor explains the root cause of overfitting. Moreover, this embedding reinitialization approach may result in the loss of substantial information, potentially resulting in suboptimal model performance. Nevertheless, it provides important inspiration for subsequent research. Another method, proposed by Wang et al. (2025), adopts an LLM-style paradigm. In this approach, generative pretraining is used to obtain embeddings for categorical features, which remain frozen during subsequent training. This strategy also helps to mitigate overfitting. However, it requires significant training resources to obtain the pretrained embeddings, and the performance cannot be fairly compared to single-epoch training because the parameter budget used during the pretraining phase is not included. Therefore, it cannot be regarded as a solution to the one-epoch problem, although it does highlight the pivotal role of the embedding layer in this phenomenon.
@@
There are several well-established and general-purpose methods to mitigate overfitting, such as dropout (Srivastava et al. (2014)), $L_{1}$ and $L_{2}$ regularization (Schmidt et al. (2009); Moore and DeNero (2011)), and weight decay (Hanson and Pratt (1988)). However, in real-world industrial applications, data is typically high-dimensional and sparse, with different features exhibiting varying degrees of sparsity. Consequently, methods like AdamW (Loshchilov and Hutter (2017)), which apply the same weight decay across all parameters, tend to be suboptimal for addressing overfitting. Specifically, such approaches may reduce the fitting accuracy for relatively dense features while failing to effectively control overfitting in sparse features.
@@ 2 Preliminary
In the industrial ASR domain, mainstream estimation models typically combine embedding layers with a variety of dense multi-layer perceptron (MLP) backbones, such as Gu et al. (2022). Taking an e-commerce platform estimation model as an example, these models process large-scale sparse inputs, including item IDs, brand IDs, and seller IDs, where the scale of categorical features can range from millions to billions. Before being processed by the estimation model, these sparse features are first mapped to low-dimensional dense representations via embedding layers. The one-hot categorical feature ${\bm{x}}_{i}(t)\in\mathbb{R}^{N_{i}}$ is defined as ${\bm{x}}_{i}(t)=\begin{bmatrix}0,0,\ldots,1,\ldots,0\end{bmatrix}^{\top},i\in[S],t\in[T]$, which can be represented by an embedding vector ${\bm{e}}_{i}\left(t\right)\in\mathbb{R}^{d_{i}}$ shown below
@@ 2.1 Neural Network Definition
where $L$ denotes the number of MLP layers, $\sigma(\cdot)$ is the ReLU function (Agarap (2018)), ${\bm{W}}_{l}$ is the linear projection matrix of layer $l,l\in[L]$. In this paper, we only discuss the impact of embedding layer so we can define the $t$-th sample output of DNN as
@@ 2.2 Rademacher Complexity Bound
Rademacher complexity is a standard tool to control the uniform convergence of given classes of predictors (Koltchinskii and Panchenko (2002),Zhang et al. (2016)). Formally, given a real-valued function class $\mathcal{H}$ with the input $\bm{z}$}, we define the empirical Rademacher complexity $\widehat{\mathcal{R}}_{T}\left(\mathcal{H}\right)$ as
@@
Based on Theorem 1 of Golowich et al. (2018), and noting that the embedding layer can be regarded as a linear projection, we can obtain the following upper bound on the Rademacher complexity of equation 3 as
@@
where $\|\cdot\|$ denotes $\ell_{2}$ norm, and each matrix ${\bm{W}}_{l}$ in function $f$ has Frobenius norm at most $M_{F}(l),l\in[L]$. $M_{E_{i}}$ is the Frobenius norm of embedding matrix ${\bm{E}}_{i}$ and the activation functions are assumed to be 1-Lipschitz and positive-homogeneous. Furthermore, because $\forall i,\|{\bm{x}}_{i}(t)\|^{2}=1$ this upper bound can be rewritten as
$\widehat{\mathcal{R}}_{T}(\mathcal{H}_{L})\leq\sqrt{\frac{S}{T}}\left(\prod_{l=1}^{L}M_{F}(l)\right)\left(\sqrt{\sum_{i=1}^{S}\sum_{j=1}^{N_{i}}\tau_{ij}}\right)\left(\sqrt{2\log(2)L}+1\right)$ (6)
where $\tau_{ij}$ is the squared $\ell_{2}$ norm of $j$-th row of embedding matrix ${\bm{E}}_{i}$. In ASR applications, the majority of parameters are concentrated in the embedding layers. Thus, we can see that $\sum_{i=1}^{S}\sum_{j=1}^{N_{i}}\tau_{ij}$ has a significant impact on the upper bound of Rademacher complexity, and consequently, on the generalization error bound. In appendix D, we provide an upper bound on the Rademacher complexity for an FM-like model, which indicates that the embedding layers also have a significantly impact on the bound.
@@ 3 The Proposed Approach
where $\varphi(\tau_{ij})=\min_{\|{\bm{e}}_{ij}\|^{2}\leq\tau_{ij}}\mathcal{L}({\bm{e}}_{ij})$ with $i\in[S]$ and $j\in[N_{i}]$. Here, ${\bm{e}}_{ij}$ denotes the $j$-th row of embedding matrix ${\bm{E}}_{i}$, and $\mathcal{L}({\bm{e}}_{ij})$ represents the average cross-entropy (CE) loss evaluated on the DNN output when the training sample activates ${\bm{e}}_{ij}$. For the CE loss function, $\mathcal{L}({\bm{e}}_{ij})$ is lower semicontinuous and bounded below on bounded sets (Goodfellow et al. (2016)). Consequently, $\varphi(\tau_{ij})$ is well-defined (Ok (2011)) and monotonically non-increasing in $\tau_{ij}$, since the feasible set expands with $\tau_{ij}$. Here, $\tau_{ij}$ is the squared norm budget of embedding vector ${\bm{e}}_{ij}$. According to DiBenedetto (2016), $\varphi(\tau_{ij})$ is differentiable almost everywhere, and we assume that $\varphi(\tau_{ij})$ is differentiable [^1] In appendix C, we give a discussion for the non-smooth case. at the optimal points $\tau^{*}_{ij}$ for analysis. The coefficient $m_{ij}$ denotes the sample frequency with which embedding vector ${\bm{e}}_{ij}$ appears in the training dataset and $C$ is a constant that defines the global norm upper bound for embedding layers. As shown in equation 6, the value of $C$ directly determines the Rademacher complexity upper bound.
@@
Proof. Based on the above assumption, at points of differentiability where a standard constraint qualification holds for the inner optimal solution, the envelope theorem and Lagrangian decomposition (Shapiro (1979)) yield
@@
Based on the KKT condition (Boyd and Vandenberghe (2004)), we have $m_{ij}\varphi^{\prime}(\tau^{*}_{ij})+\mu_{0}-\mu_{ij}=0$. Applying the complementary slackness condition, we obtain $\mu_{ij}\tau^{*}_{ij}=0$. This implies the optimality condition for $\tau_{ij}>0$ can be simplified as
@@ 3.1 Adaptive Regularization Method
As shown in equation 6, embedding layers in ASR applications typically take the majority of the model parameters and have a substantial impact on the generalization error bound. Proposition 1 suggests that the norm budget for each embedding vector should be allocated according to its sample frequency. However, it is not easy to use the frequency directly during the training process. We can estimate $m_{ij}$ via the stochastic occurrence interval $\mathcal{I}_{ij}$ of ${\bm{e}}_{ij}$ for $i\in[S]$ and $j\in[N_{i}]$. Specifically, given that ${\bm{x}}_{i}(t)$ is sampled from an i.i.d. distribution (a common assumption in deep learning), we have $\mathbb{E}[m_{ij}]=T/\mathbb{E}[\mathcal{I}_{ij}]$ (Papoulis (1965)). It enables us to incorporate frequency estimation into the norm budget allocation strategy during training.
Based on the analysis above, we propose an adaptive method that assigns the regularization strength based on the occurrence interval of each embedding vector. Specifically, at each training step $k$, we define last valid update step (LVS) for the embedding vector ${\bm{e}}_{ij}$ as $s^{k}_{ij}$. If the gradient norm of ${\bm{e}}_{ij}$ satisfies $\|{\bm{g}}_{ij}\|>0$, we update the LVS by setting $s^{k}_{ij}=k$. Otherwise, $s^{k}_{ij}$ retains its previous value. Therefore, $s^{k}_{ij}$ serves as a lazy-update variable. We define the update interval of step $k$ as $I^{k}_{ij}=k-s^{k-1}_{ij}-1$. The adaptive regularization coefficient $\lambda^{k}_{ij}$ is then dynamically computed as below
$\lambda^{k}_{ij}=\min\left(1,\alpha I^{k}_{ij}\right),i\in[S],j\in[N_{i}]$ (11)
Here, $\alpha\in[0,1)$ denotes the base regularization coefficient. Following the decoupled weight decay approach in AdamW (Loshchilov and Hutter (2017)), we incorporate the dynamically computed regularization into each optimizer update step. Algorithm 1 outlines the procedure of Adam with adaptive regularization (AdamAR).
@@
In fact, our method is not only suitable for Adam, but also compatible with various gradient-based optimizers which have weight decay factor, such as Adagrad (Duchi et al. (2011)). We have implemented an adaptive version for Adagrad in algorithm 2 of the appendix G. In section 4, we conducted comparative experiments evaluating performance across both Adam and Adagrad.
Moreover, the AdamAR algorithm requires additional storage to record the last valid update step. Further details of the computation and memory analysis are provided in appendix F.
@@ 3.2 Discussion on the Mechanism of Regularization
When adaptive regularization is applied according to equation 11, the update rule for parameters satisfy $\|\theta^{k}_{p}\|\leq\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|+\|\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}+\varepsilon})\|$
@@
$\|\theta^{k}_{p}\|\leq\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|+\|\eta\cdot\hat{m}^{k}_{p}/(\sqrt{\hat{v}^{k}_{p}+\varepsilon})\|$ (13)
From proposition 2, the update rule for $\theta^{k}_{p}$ can be interpreted intuitively. If the interval $I^{k}_{p}$ is large for the corresponding sparse categorical features, the term $\left(1-\alpha\right)^{I^{k}_{p}}\|\theta^{k-1}_{p}\|$ becomes negligible, and $\theta^{k}_{p}$ is effectively determined by the latest gradient. In the ASR domain, certain categorical feature values may not appear in every mini-batch, resulting in their embedding parameters being updated infrequently, while the MLP parameters are updated in every batch. After the MLP component has nearly converged, the embedding parameters associated with sparse features will have received far fewer updates and may become misaligned with the current state of the MLP parameters, making the previous value $\theta^{k-1}_{p}$ less informative. Proposition 2 demonstrates that the factor $\left(1-\alpha\right)^{I^{k}_{p}}$ exponentially attenuates the previous value $\theta^{k-1}_{p}$, making the embedding parameters of sparse features depend more heavily on the current gradient. Meanwhile, since the update interval $I^{k}_{p}$ for MLP parameters is identically zero (as they are updated in every batch), the regularization primarily affects the embedding parameters corresponding to very low-frequency feature values. Moreover, we observe that the method proposed in Liu et al. (2023) is a special case of our approach when zero reinitialization is applied to the embedding and $I^{k}_{p}$ is specified as
@@
3.3 Minimum Convergence
We analyze the convergence of adaptive regularization methods in the non-convex setting using the minimum convergence rules proposed by Khaled and Richtárik (2020). The assumptions underlying our convergence analysis are listed below.
Assumptions 1. The function $f$ is differentiable and its gradient is Lipschitz continuous, i.e., there exists $L_{0}>0$ such that $\|\nabla f(\bm{\theta}^{k+1})-\nabla f(\bm{\theta}^{k})\|\leq L_{0}\|\bm{\theta}^{k+1}-\bm{\theta}^{k}\|,\forall k\geq 1$, and $f$ is lower bounded at the optimal solution, i.e., $f^{*}>$-$\infty$.
Assumptions 2. $\bm{g}^{k}$ is an unbiased estimator of the full gradient, i.e., $\mathbb{E}\left[\bm{g}^{k}\right]=\nabla f(\bm{\theta}^{k})$ with $M>0$, and the algorithm accesses a bounded stochastic gradient, i.e., $\|\bm{g}^{k}\|\leq M$ a.s.
Proposition 3
The adaptive regularization method preserves the minimum convergence bound of the Adam optimizer with stochastic conditions, which can be expressed as
$\min_{1\leq k\leq K}\mathbb{E}\left[\left\|\nabla f\left(\bm{\theta}^{k}\right)\right\|^{2}\right]\leq\frac{C_{1}+C_{2}\sum_{k=1}^{K}\eta_{k}+C_{3}\sum_{k=1}^{K}\eta_{k}^{2}}{\sum_{k=1}^{K}\eta_{k}}$ (15)
where $C_{1}$, $C_{2}$ and $C_{3}$ are constants, $\eta_{k}$ denotes the step size at iteration $k$ out of $K$ total iterations.
The proof, along with the definitions of the constants $C_{1}$, $C_{2}$, and $C_{3}$, can be found in appendix B.1. Proposition 3 shows that neither the random noise nor the deterministic regularization parameter affects the minimum convergence of Adam, and it only changes the constant term in equation 15.